(7%) Problem 4: A 66 kg and a 82 kg skydiver jump from an airplane at an altitude of 6000 m, both falling in the diving/headfirst position. Assume their surface area is 0.105 m and the drag coefficient is 0.70 Randomized Variables ,-66 kg m 82kg A-0105 m 50% Part (a) How long will it take for the first dner to reach the pound in seconds assuming the time to reach term na raloct Grade Summary Potential95 Submissio s man and the density of the air is a constant 1.21 kg/m)7 tan0 5 6 Attempts remaning atan0 cosh tan cotanh0 500 * Degrees Radians SbeH ie up deductior par feedback Hint Hists --deduction per hunt Hints remainin Feedback: ム50% Part (b) Hor long wa t take r the second s di er to reach the ground m seconds (assuming the time to reach terminal velocity is small)?
Answer #1

assuming time to reach terminal velocity is small then we ca consider that diver travels whole distance with terminal velocity.

at terminal velocity, vt : F_net = F_drag - Fg = 0

Cd rho A v^2 /2 = m g

(A) (0.70)(1.21)(0.105) v^2 /2 = 66 x 9.81

v = 120.6 m/s

t = 6000 / 120.6 = 50 sec  

(B) (0.70)(0.105)(1.21) v^2 /2 = 82 x 9.8

v = 134.4 m/s

t =45 sec